# Multi-tracer beyond linear theory

### v1

@@ Appendix B Dependence on the fiducial values
In this section, we discuss how our results depend on the fiducial values chosen for the bias parameters in Eq. (2.26). The right panel of Fig. 11 shows the ratio between the MT and ST error bars as a function of the fiducial values for $b_{1}$, $b_{\delta^{2}}$, and $b_{\mathcal{G}_{2}}$. In this analysis, we make all the fiducial bias parameters change together. For the MT results, we considered splits of $\Delta b=0.3$ (dashed) and $\Delta b=0.6$ (solid). We observe fluctuations of approximately $20\%$ in the curves, yet the overall qualitative results remain unchanged.
@@ Appendix C Extra plots for $b_{\rm all}^{+\pm\pm\pm}$ and different $k_{\rm max}$
In this Appendix, we include complementary plots that were not added to the main part of the paper. In Fig. 12, the left panel displays all possible combinations of $b_{\rm all}^{+\pm\pm\pm}$; although we observe some relative differences between them, no qualitative change occurs. In the right panel, we present the same plot as in Fig. 1, but with $k_{\rm max}=0.05\,h/$Mpc. The error bars are, as expected, larger, but we find that a split in $b_{1}$ becomes relatively more important than a split in $b_{\delta^{2}}$.

### v2

@@ Appendix B Dependence on the fiducial values
In this section, we discuss how our results depend on the fiducial values chosen for the bias parameters in Eq. (2.26). The left panel of Fig. 13 shows the ratio between the MT and ST error bars as a function of the fiducial values for $b_{1}$. The right panel of Fig. 13 shows the ratio between the MT and ST error bars as a function of the fiducial values for $b_{\delta^{2}}$, $b_{\mathcal{G}_{2}}$ and $b_{\Gamma_{3}}$, changing all those fiducial bias parameters together. For the MT results, we considered splits of $\Delta b=0.3$ (dashed) and $\Delta b=0.6$ (solid). We observe very small fluctuations when changing the fiducial value of $b_{1}$ and fluctuations of at most $20\%$ when varying the fiducial value of the non-linear bias parameters, yet the overall qualitative results remain unchanged.
Figure 13: Ratio between MT and ST error bars as a function of the fiducial bias parameters chosen in Eq. (2.26). For the MT results, we consider $\Delta b=0.3$ (dashed) and $\Delta b=0.6$ (solid) splits. In the left, we vary $b_{1}^{\rm fid}$ keeping the other parameters fixed to Eq. (2.26), in the right we vary the non-linear bias fiducial values, keeping $b_{1}^{\rm fid}=1.16$.
@@ Appendix C Extra plots for $b_{\rm all}^{+\pm\pm\pm}$ and different $k_{\rm max}$
In this Appendix, we include complementary plots that were not added to the main part of the paper. In Fig. 14, the left panel displays all possible combinations of $b_{\rm all}^{+\pm\pm\pm}$; although we observe some relative differences between them, no qualitative change occurs. In the right panel, we present the same plot as in Fig. 1, but with $k_{\rm max}=0.05\,h/$Mpc. The error bars are, as expected, larger, but we find that a split in $b_{1}$ becomes relatively more important than a split in $b_{\delta^{2}}$. We also show in Fig. 15 the error in the cosmological parameters for a split $\Delta b_{\delta^{2}}$ (solid) and $\Delta b_{\mathcal{G}_{2}}$ (dashed) normalized by the error calculated using $\Delta b_{1}$ as a function of $k_{\rm max}$. As we restric the analysis to large scales, the split in the linear bias parameter becomes more relevant than a split in the non-linear bias.
Figure 15: Error in $\omega_{\rm cdm}$ (top), $h$ (middle) and $A_{s}$ (bottom) for a split $\Delta b_{\delta^{2}}$ (solid) and $\Delta b_{\mathcal{G}_{2}}$ (dashed) divided by the error calculated using a split in $\Delta b_{1}$ as a function of $k_{\rm max}$. Different colors represent different $\Delta b$ values. We see that as we consider only linear scales ($k_{\rm max}\to 0$), a split in $b_{1}$ tends to be more efficient than a split in non-linear bias parameters.